40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 26 21:01:26 2019, gautam, Update, LSC, POX and POY locking 
    Reply  Thu Jun 27 14:28:12 2019, gautam, Update, LSC, POX and POY locking 
       Reply  Tue Jul 2 17:34:54 2019, gautam, Update, LSC, Irides opened up on EY table 
          Reply  Tue Jul 9 18:06:05 2019, gautam, Update, LSC, Y-arm ASS in a workable state ditherFreqs.pdftransRenorm.png
Message ID: 14705     Entry time: Thu Jun 27 14:28:12 2019     In reply to: 14704     Reply to this: 14720
Author: gautam 
Type: Update 
Category: LSC 
Subject: POX and POY locking 

After a more systematic alignment effort, I was able to get the spots better centered on the optics (judged by eye from the analog camera views). TRY ~0.7, TRX~1.15. The X-arm dither alignment system seems to work out-of-the-box with the existing settings, I was able to run it and maximize the X-arm transmission.

Other work: I also cleaned up the area around MC2 a litte - laptop from on top of the vacuum chamber was removed and a rogue ethernet cable was also removed. The resulted in some misalignment of the IMC, which I corrected by manual alignment. Now the IMC is locked again with nominal transmission levels.

On the PSL table, I re-routed the RF output from the BeatMouth to the regular IR-ALS electronics chain (it was hijacked for PLL investigations). At EX, I disconnected the cable running from the LB1005 to the EX NPRO laser PZT (again was being used for PLL locking), and re-connected the output from the Green uPDH box to allow for some ALS tests to be done. I could then lock the EX green beam to the X-arm, and achieved GTRY ~ 0.35 using the ASX system. More to follow on ALS tests later today.

ELOG V3.1.3-